how do i plot d=240-15t on a graph?
Answer by TimothyLamb(4379)

You need two points on the line in order to graph it, but first get the equation in slope-intercept form: y = mx + b
---
For this problem, t (normally x) is plotted along the horizontal axis, and d (normally y) is plotted along the vertical axis.
---
d = 240 - 15t
d = -15t + 240
---
Plug-in two different t values in the equation to get their d values, like so:
---
Point p1, use t=0:
d = -15(0) + 240
d = 240
p1 = (0, 240)
---
Point p2, use t=1:
d = -15(1) + 240
d = 225
p2 = (1, 225)
---
Now draw a line between points p1 and p2 to graph the equation.
The line extends infinitely beyond both points.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php